OpenStudy (anonymous):

How do you calculate the area of a circle knowing the diameter and or the radius?

OpenStudy (angela210793):

A=pi*r^2

OpenStudy (anonymous):

A=pi*r^2 A=pi*(d/2)^2

OpenStudy (angela210793):

r=d/2

OpenStudy (amistre64):

\[2pi\int_{0}^{x}r\ dr\]

OpenStudy (anonymous):

Thanks to all!!

OpenStudy (anonymous):

To find the area, square the radius and then multiply by pi (~3.14). For example, if the radius is 3, then 3*3 = 9, so 9*pi or ~28.27